# Do Teachers Dream of GenAI Widening Educational (In)equality? Envisioning the Future of K-12 GenAI Education from Global Teachers’ Perspectives

### v3

@@ 1. Introduction
Today, generative artificial intelligence (GenAI) is rapidly entering K–12 classrooms across the globe (Mintz et al., 2023; Mathur, 2025). Whether teachers welcome it or not, students are already engaging with GenAI tools in and beyond the classroom (Adisa and Adefisayo, 2025), for instance, using ChatGPT to draft essays (Levine et al., 2025), solve math problems (Walkington et al., 2025), or generate summaries of readings (Steele, 2023). This rapid spread makes GenAI not only a technological trend but also an urgent equality issue at the societal level. On one hand, it has the potential to democratize learning opportunities, providing personalized tutoring and cognitive support that could help close long-standing gaps in access and achievement (Tafazoli, 2024; James and Andrews, 2024). On the other hand, it risks exacerbating divides by disproportionately benefiting students with reliable infrastructure, digital literacy, and supportive school resources, while leaving behind those in under-resourced contexts (Hendawy, 2024; Murgia et al., 2024).
@@ 3.1. Context: AI Education in the United States, South Africa, and Taiwan
Taiwan combines advanced high-tech development with cultural and linguistic particularities. Under the centralized “108” Curriculum Guidelines (108課綱) (Ministry of Education, Taiwan, 2025)), IT (including AI literacy) is mandatory from middle school, supported by nationwide teacher-training platforms like the Adaptive Learning Platform (ADL, or 因材網) (Ministry of Education, Taiwan, 2025; Lee, 2024). The government funds devices and resources for rural and indigenous schools to reduce disparities (National Development Council, Taiwan, 2025). However, differences between Traditional Chinese and Simplified Chinese corpora pose challenges for localizing large language models, given the larger scale and dominance of Simplified corpora compared to the smaller, Taiwan-focused Traditional corpora (Lyu et al., 2025).
@@ 4.1. Using GenAI Education to Alleviate Pre-Existing Educational Inequalities
Teachers consistently described that GenAI education can be used to mitigate inequalities that had long predated GenAI’s arrival in classrooms. Such inequalities, including unequal access to knowledge and opportunities, had historically taken multiple forms: some students lacked basic awareness of emerging technologies (subsubsection 4.1.1); others struggled with shortages of learning resources (subsubsection 4.1.2); many came to school with weaker digital literacy skills that excluded them from computing participation (subsubsection 4.1.3); and marginalized learners, such as those with disabilities or without access to advanced courses, were often denied pathways to develop essential competencies (subsubsection 4.1.4).
@@ 4.2. Preventing GenAI from Creating New Inequalities Through Teaching Practice
T10 in Taiwan illustrated this challenge through a simple case of linguistic variation. She noticed that GenAI often defaulted to Mandarin Chinese terms common in mainland China, rather than the expressions used in Taiwan. To address this, she deliberately turned such discrepancies into class activities: she showed students examples, asked them to identify differences, and facilitated discussions on why the AI made these choices. “In mainland China they say Bo-Luo (菠萝), in Taiwan we say Feng-Li (鳳梨)… I use this to make students aware of local linguistic differences in AI outputs.” By explicitly drawing attention to these differences, T10 encouraged her students to critically examine AI outputs, validate their own linguistic identity, and understand that cultural variation is a form of richness often ignored by current GenAI systems.
@@ 5. Findings: Continuing Inequalities Challenges in GenAI Education Beyond Individual Efforts (RQ2)
Infrastructural Constraints on GenAI (1) Increased device availability without reliable connectivity.
Education (2) Insufficient basic digital literacy among students with limited access to technology.
@@
Insufficient Training and Curricular (1) Uneven access to high-quality professional development for GenAI instruction.
Guidance (2) A widening gap between GenAI development and educators’ instructional preparedness.
@@ 7. Discussion
This study contributes to a growing understanding of how K–12 teachers navigate the promises and perils of GenAI in educational equality through their everyday GenAI education teaching practices. Our findings across three regions (the United States, South Africa, and Taiwan) highlight a threefold dynamic. First, teachers actively used GenAI education as a mechanism to alleviate pre-existing inequalities and to prevent the emergence of new ones (section 4). Second, despite these efforts, their work was repeatedly constrained by infrastructural gaps, limited training, and restrictive social norms that lay beyond individual control (section 5). Third, teachers further envisioned concrete pathways for the future of GenAI education, outlining guidelines for schools, companies, and governments to support equal access and meaningful use better (section 6).
In the following discussion section, first, we reconceptualize GenAI education as a teacher practice of negotiating equality (see subsection 7.1). Second, we distinguish GenAI education from earlier forms of computing education, highlighting how its natural language interface lowers barriers for diverse learners but also introduces new equality challenges that demand teacher mediation (see subsection 7.2). Third, we extend beyond prior studies on individual adoption to demonstrate that teachers’ equality-oriented practices are deeply constrained by structural conditions of infrastructure, training, and social norms, underscoring the need to bridge micro-level pedagogical work with macro-level systemic inequalities (subsection 7.3). Finally, we synthesize these findings into a set of design tensions that articulate broader HCI contributions, and derive design and policy implications for supporting equality-oriented GenAI education at scale (subsection 7.4, subsection 7.5).
@@ 7.4. Design Tensions in Equality-Oriented GenAI Education
For HCI, this tension extends beyond education by foregrounding personalization as a multi-level phenomenon. While systems may successfully tailor content to individual users, they can simultaneously standardize values and representations at scale (Li et al., 2025; Xiao et al., 2025d; Schecter and Richardson, 2025). This insight urges HCI researchers to examine not only who personalization serves, but also whose cultural and epistemic assumptions are embedded and amplified through personalized AI systems.
@@ 7.5. Design and Policy Implications
Several of the tensions identified in Section 7.4—particularly those between lowering barriers and deepening new inequalities, personalization and standardization of values, and human care versus machine-mediated support—directly implicate the design of GenAI systems.
@@
Finally, addressing the tension between human care and machine-mediated support requires design to move closer to pedagogy. GenAI systems for educational contexts should embed features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to engage with it ethically and critically (Xiao et al., 2025d). Teachers’ negative experiences also highlight the need for stronger AI literacy among developers, managers, and decision-makers, so that potential harms, especially those affecting vulnerable learners, can be anticipated rather than treated as afterthoughts, as emphasized by prior CHI research (Xie et al., 2025).
@@
Teacher capacity likewise requires systemic support. Professional development must go beyond tool introduction toward sustained pedagogical frameworks for responsible GenAI integration, enabling teachers to manage the tensions between access, critical engagement, and ethical use over time (Ding et al., 2024). Our findings also underscore the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped students’ access to GenAI, suggesting the need for public education and participatory approaches to policy-making, consistent with HCI and CSCW work on community-centered design (Villari and others, 2021; Zytko et al., 2022).

### v4

@@ 1. Introduction
Today, generative artificial intelligence (GenAI) is rapidly entering K–12 classrooms across the globe (Mintz et al., 2023; Mathur, 2025; Xiao et al., 2025d). Whether teachers welcome it or not, students are already engaging with GenAI tools in and beyond the classroom (Adisa and Adefisayo, 2025), for instance, using ChatGPT to draft essays (Levine et al., 2025), solve math problems (Walkington et al., 2025), or generate summaries of readings (Steele, 2023). This rapid spread makes GenAI not only a technological trend but also an urgent equality issue at the societal level. On one hand, it has the potential to democratize learning opportunities, providing personalized tutoring and cognitive support that could help close long-standing gaps in access and achievement (Tafazoli, 2024; James and Andrews, 2024). On the other hand, it risks exacerbating divides by disproportionately benefiting students with reliable infrastructure, digital literacy, and supportive school resources, while leaving behind those in under-resourced contexts (Hendawy, 2024; Murgia et al., 2024).
@@ 3.1. Context: AI Education in the United States, South Africa, and Taiwan
Taiwan combines advanced high-tech development with cultural and linguistic particularities. Under the centralized “108” Curriculum Guidelines, IT (including AI literacy) is mandatory from middle school, supported by nationwide teacher-training platforms like the Adaptive Learning Platform (ADL) (Ministry of Education, Taiwan, 2025; Lee, 2024). The government funds devices and resources for rural and indigenous schools to reduce disparities (National Development Council, Taiwan, 2025). However, differences between Traditional Chinese and Simplified Chinese corpora pose challenges for localizing large language models, given the larger scale and dominance of Simplified corpora compared to the smaller, Taiwan-focused Traditional corpora (Lyu et al., 2025).
@@ 4.1. Using GenAI Education to Alleviate Pre-Existing Educational Inequalities
Teachers consistently described that GenAI education can be used to mitigate inequalities that had long predated GenAI’s arrival in classrooms. Such inequalities, including unequal access to knowledge and opportunities, had historically taken multiple forms: some students lacked basic awareness of emerging technologies (see subsubsection 4.1.1); others struggled with shortages of learning resources (see subsubsection 4.1.2); many came to school with weaker digital literacy skills that excluded them from computing participation (see subsubsection 4.1.3); and marginalized learners, such as those with disabilities or without access to advanced courses, were often denied pathways to develop essential competencies (see subsubsection 4.1.4).
@@ 4.2. Preventing GenAI from Creating New Inequalities Through Teaching Practice
T10 in Taiwan illustrated this challenge through a simple case of linguistic variation. She noticed that GenAI often defaulted to Mandarin Chinese terms common in mainland China, rather than the expressions used in Taiwan. To address this, she deliberately turned such discrepancies into class activities: she showed students examples, asked them to identify differences, and facilitated discussions on why the AI made these choices. “In mainland China they say Bo-Luo, in Taiwan we say Feng-Li… I use this to make students aware of local linguistic differences in AI outputs.” By explicitly drawing attention to these differences, T10 encouraged her students to critically examine AI outputs, validate their own linguistic identity, and understand that cultural variation is a form of richness often ignored by current GenAI systems.
@@ 5. Findings: Continuing Inequalities Challenges in GenAI Education Beyond Individual Efforts (RQ2)
Infrastructural Constraints on GenAI Education (1) Increased device availability without reliable connectivity.
(2) Insufficient basic digital literacy among students with limited access to technology.
@@
Insufficient Training and Curricular Guidance (1) Uneven access to high-quality professional development for GenAI instruction.
(2) A widening gap between GenAI development and educators’ instructional preparedness.
@@ 7. Discussion
This study contributes to a growing understanding of how K–12 teachers navigate the promises and perils of GenAI in educational equality through their everyday GenAI education teaching practices. Our findings across three regions (the United States, South Africa, and Taiwan) highlight a threefold dynamic. First, teachers actively used GenAI education as a mechanism to alleviate pre-existing inequalities and to prevent the emergence of new ones (see section 4). Second, despite these efforts, their work was repeatedly constrained by infrastructural gaps, limited training, and restrictive social norms that lay beyond individual control (see section 5). Third, teachers further envisioned concrete pathways for the future of GenAI education, outlining guidelines for schools, companies, and governments to support equal access and meaningful use better (see section 6).
In the following discussion section, first, we reconceptualize GenAI education as a teacher practice of negotiating equality (see subsection 7.1). Second, we distinguish GenAI education from earlier forms of computing education, highlighting how its natural language interface lowers barriers for diverse learners but also introduces new equality challenges that demand teacher mediation (see subsection 7.2). Third, we extend beyond prior studies on individual adoption to demonstrate that teachers’ equality-oriented practices are deeply constrained by structural conditions of infrastructure, training, and social norms, underscoring the need to bridge micro-level pedagogical work with macro-level systemic inequalities (see subsection 7.3). Finally, we synthesize these findings into a set of design tensions that articulate broader HCI contributions, and derive design and policy implications for supporting equality-oriented GenAI education at scale (see subsection 7.4, subsection 7.5).
@@ 7.4. Design Tensions in Equality-Oriented GenAI Education
For HCI, this tension extends beyond education by foregrounding personalization as a multi-level phenomenon. While systems may successfully tailor content to individual users, they can simultaneously standardize values and representations at scale (Li et al., 2025; Xiao et al., 2025e; Schecter and Richardson, 2025). This insight urges HCI researchers to examine not only who personalization serves, but also whose cultural and epistemic assumptions are embedded and amplified through personalized AI systems.
@@ 7.5. Design and Policy Implications
Several of the tensions identified in subsection 7.4—particularly those between lowering barriers and deepening new inequalities, personalization and standardization of values, and human care versus machine-mediated support—directly implicate the design of GenAI systems.
@@
Finally, addressing the tension between human care and machine-mediated support requires design to move closer to pedagogy. GenAI systems for educational contexts should embed features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to engage with it ethically and critically (Xiao et al., 2025e). Teachers’ negative experiences also highlight the need for stronger AI literacy among developers, managers, and decision-makers, so that potential harms, especially those affecting vulnerable learners, can be anticipated rather than treated as afterthoughts, as emphasized by prior CHI research (Xie et al., 2025).
@@
Teacher capacity likewise requires systemic support. Professional development and corresponding assessments (Zhang et al., 2026) must go beyond tool introduction toward sustained pedagogical frameworks for responsible GenAI integration, enabling teachers to manage the tensions between access, critical engagement, and ethical use over time (Ding et al., 2024). Our findings also underscore the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped students’ access to GenAI, suggesting the need for public education and participatory approaches to policy-making, consistent with HCI and CSCW work on community-centered design (Villari and others, 2021; Zytko et al., 2022).